OpenStudy (anonymous):

Subtract. (9x8+ 4x7 + 7x6 – 2) – (7x8 + 3x7 + 7x6 – 2)

OpenStudy (anonymous):

we can only add and subtract like terms

OpenStudy (anonymous):

\[\LARGE (+)\cdot (-)=(-)\] \[\LARGE (-)\cdot (+)=(-)\] \[\LARGE (-)\cdot (-)=(+)\] \[\LARGE (+)\cdot (+)=(+)\]

OpenStudy (anonymous):

9x^8 + 4x^7 + 7x^6 - 2 - 7x^8 - 3x^7 - 7x^6 +2 =2x^8 + x^7

OpenStudy (campbell_st):

collect the like terms \[(9 - 7)x^2 +(4 - 3)x^7 +(7 - 7)x^6 +(-2 + 2) = 2x^8 + x^7\]